Origin of the grid Some text with the bottom left (i.e. “bl” exactly where one wants it This text was created using a parbox.

• Multiple line paragraphs go inside parboxes

• The top left (tl) position is controlled by pstricks

• You can control various itemize position features but that requires the enumitem package.

• Text can run off the end of the fbox, you need to control the height of the parbox if you want it inside the fbox.

• Note that the position of this parbox is controlled by pstricks even though it is not inside the displayed grid.

This is a hyperlink to Tom Schneider’s L^{A}TEXpage version = 1.03 of griddemo.tex 2015 May 25 Claude Shannon wrote this equation: H = − ∑_{i}^{M} p_{i} log_{2} p_{i} (bits per symbol)

Demonstration of precise page placement using PSTricks

Demo of reading in a postscript program or eps: